OpenStudy (anonymous):

1 -- -4 x^2 --------- 4 4 -- + -- +1 x^2 x

OpenStudy (anonymous):

just need it simplify

jhonyy9 (jhonyy9):

(1-x2)/x2 * x2/(4+4x +x2) =(1-x2)/(x2+4x+4) =(1-x2)/(x+2)2 (x+2)2 is (x+2) squared

OpenStudy (anonymous):

so my final is 1-x^2 ------ (x+2)^2

OpenStudy (phi):

If the numerator is \[ \frac{1}{x^2} -4\] then you get a different answer. The easiest way to simplify is to multiply the top and bottom by x^2 the top becomes 1-4x^2 the bottom becomes 4+4x+x^2 As you know, the top can be factored into (1-2x)(1+2x) and the bottom into (x+2)^2